I need some assistance on the whole problem and how to do it please. How do I do the set up and then solving it along with check. Please Please help.
The area of a triangle is given as A = 1/2 BH. Find the height (H) given that the Base (B) = 14 inches and the area (A) = 56 sq. inches. I need to find the height.

Answer by checkley71(8403)

A=1/2BH
56=1/2*14*B
56=7*B
B=56/7
B=8 INCHES FOR THE BASE